OpenStudy (anonymous):

Missy is organizing the seating arrangements for a concert in her local park. The first row's seats are modeled by the function f(x) = 10. To give everyone a good view, each row must have two more seats than the row before it. The number of additional seats a row has is modeled by the function s(x) = 2(x - 1). Explain to Missy how to create an equation to predict the number of seats in any row. Describe how to determine the number of seats in the 20th row. Use complete sentences.

OpenStudy (here_to_help15):

Your overall function will be 2(x-1) + 10. Do you see when you plug in 1 for x, you will get 10 as answer? x=1 represents the first row. x=20 will represent the 20th row.

OpenStudy (anonymous):

im soo confused lol i understand how we got the overall function but how would i "explain that to missy"

OpenStudy (here_to_help15):

lol slapped the sense into "missy"

OpenStudy (anonymous):

could you help me think of a way to write that out in a sentence

OpenStudy (here_to_help15):

I can try lol

OpenStudy (anonymous):

thanks you!!

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (here_to_help15):

Normally, the # seats in row #n would be [ 10 + 2(n-1) ]. Indeed, #1 has 10 + 2(1-1) = 10 seats #2 has 10 + 2(2-1) = 12 seats, etc... But you now specify that the # of *additional* seats is s(x) = 2(x-1), or, s(#n) = 2(#n-1) additional seats. So in all, row #n would have [ 10 + 2(n-1) ] + s(#n) = 10 + 2(n-1) + 2(n-1) = 4n + 6.

OpenStudy (anonymous):

okay thank you!